Calculate the concentration of OH- in a 0.15 M solution of NH3. Kb = 1.8 × 10-5

1
Expert's answer
2022-05-23T06:48:01-0400

NH3 + H2O = NH4+ + OH-


1) Kb = [NH4+] * [OH-] / [NH3] .where, [NH4+] and [OH-] are equal, so we denote them by X

Kb = X * X / [NH3] or

Kb = X2 / [NH3]


2) 1,8 * 10-5 = X2 / 0,15

X=√(1,8*10-5 * 0,15) =√2,7*10-6 = 1,64 *10-3

ANSWER: [ OH- ] = 1,64 *10-3 M